# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v3

@@ 6.4 Operational Definition of Fatigue Onset (Ground Truth)
Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$. In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates). To stabilise the series prior to defining fatigue onset, we use a simple Beta–Binomial shrinkage estimator:
@@
where $(\alpha,\beta)$ represent a weakly informative prior over the underlying click probability. We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$. In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives. We then define the smoothed CTR series as a trailing moving average of length $w_{s}$:
@@
with $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
@@
Fatigue is defined relative to a “recent best” baseline. For each day $t$, define
@@
We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive days. Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
$Drop_{t^{\prime}}\geq\delta\quad\text{for all}\quad t^{\prime}\in\{t,t+1,\dots,t+p-1\}.$
In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ days. Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
All quantities are evaluated only when the required history is available. For creatives with missing days, lookback windows are defined over the last $k$ observed days rather than calendar days. This onset definition is retrospective because it requires observing $p$ consecutive days; in an online setting, the earliest actionable alert corresponding to this definition would occur at day t+p1t+p-1.
@@
To assess sensitivity to the operational definition, we report robustness results over $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed. This ensures that comparative conclusions are not artefacts of a single threshold choice.
@@ 6.5 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to five synthetic datasets designed to mimic common real-world scenarios of creative performance decay. The resulting visualisations, shown in Figures 1 through 5, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
@@
Figure 6: Non-Continuous Data (Sharp Drop): Demonstrates the method’s robustness to significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method correctly identifies the change point despite irregular sampling.
These visualisations demonstrate the algorithm’s ability to adapt to different fatigue patterns. For a marketing analyst, these plots provide a clear and immediate narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 specifically illustrate that signature-based methods can be applied to irregularly sampled trajectories. Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
@@ 7 Results
We evaluate detection performance against the operational fatigue onset definition in Section 6.4. The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$). For each method, we define an alert date $\hat{\tau}$ as the first alert date for that creative. For the signature-based approach, $\hat{\tau}$ is the start date of the first segment classified as “declining” (with a fallback to the first detected change point). For baselines, $\hat{\tau}$ is the first day on which their heuristic trigger persists for three consecutive observed days.
@@
Table 1: Lead time and alert burden under an operational ground truth definition
Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median lead
@@ 7.1 Interpretation: early warning versus alert burden
Table 1 highlights a practical trade-off. Both the signature-based detector and the moving-average crossover provide early warning for all fatigue creatives under the operational definition, but they also generate alerts for all non-fatigue creatives in this dataset. By contrast, the rolling regression baseline is conservative (low burden) but misses most fatigue onsets and is, on average, late when it does trigger. Large positive lead times should be interpreted cautiously: extremely early alerts can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection.

### v4

@@ 6.4 Operational Definition of Fatigue Onset
Let $I_{t}$ and $C_{t}$ denote daily impressions and clicks for a given creative, and define the observed CTR as $CTR_{t}=C_{t}/I_{t}$ when $I_{t}>0$; zero-impression days are excluded from raw CTR calculations. In finite samples, $CTR_{t}$ exhibits non-trivial measurement noise (particularly when $I_{t}$ fluctuates). To stabilise the series prior to defining fatigue onset, we use a simple Beta–Binomial shrinkage estimator:
@@
where $(\alpha,\beta)$ represent a weakly informative prior over the underlying click probability. We parameterise the prior by a notional prior sample size $m_{0}$ and a global mean CTR $\bar{p}$, setting $\alpha=m_{0}\bar{p}$ and $\beta=m_{0}(1-\bar{p})$. In our experiments, we use $m_{0}=1000$ impressions and estimate $\bar{p}$ as the pooled mean CTR across all creatives, assuming $0<\bar{p}<1$. This pooled value is an offline evaluation device and means that one creative’s smoothed series depends weakly on the panel-level prior. An online deployment would need to fix $\bar{p}$ from a historical training baseline before scoring future days, or use a leave-one-creative-out or externally specified prior. We then define the smoothed CTR series as a trailing moving average of length $w_{s}$. For $t\geq w_{s}-1$,
@@
and for initial times $t<w_{s}-1$ we use the available prefix,
$\widetilde{CTR}_{t}=\frac{1}{t+1}\sum_{j=0}^{t}\widehat{p}_{t-j}.$
We set $w_{s}=7$ to mitigate weekly noise while preserving responsiveness.
@@
Fatigue is defined relative to a “recent best” baseline. Assuming $1\leq w_{s}\leq B$, for each eligible day $t$ define
@@
The drop ratio is evaluated only when the baseline window is nonempty and $Base_{t}>0$. We declare that a creative has entered fatigue when the degradation exceeds a material threshold $\delta$ for a sustained period of $p$ consecutive observed days. Formally, the fatigue onset date $\tau$ is the earliest $t$ such that
$Drop_{s_{j+r}}\geq\delta\quad\text{for all}\quad r=0,\dots,p-1,$
where $s_{j}$ denotes the ordered observed days and $\tau=s_{j}$. In the main specification, we set $\delta=0.15$ (a 15% sustained decline relative to recent-best performance) and $p=5$ observed days. The rule is based on observed-record persistence rather than calendar-day contiguity. Because the synthetic panel is complete, this distinction has no effect in the reported experiment. For sparse synthetic extensions, the rule should add a maximum-span constraint such as $s_{j+p-1}-s_{j}\leq G$ calendar days before declaring a sustained run. Creatives that never satisfy this criterion are treated as having no fatigue onset within the observed horizon.
All quantities are evaluated only when the required history is available. For creatives with missing days, let $s_{1}<\cdots<s_{n}$ be the observed days; then for $t=s_{j}$, compute $Base_{t}=\max\{\widetilde{CTR}_{s_{j-r}}:r=w_{s},\dots,B\}$ whenever those observed-day indices exist. This onset definition is retrospective because it requires observing $p$ consecutive qualifying observations; in an online setting, the earliest actionable alert corresponding to this definition would occur on the $p$-th qualifying observed day.
@@
To assess sensitivity to the operational definition, the intended robustness check varies $\delta\in\{0.10,0.15,0.20\}$ and $p\in\{3,5,7\}$, holding the smoothing parameters fixed. The present draft reports the main operating point and detector-parameter sensitivity below; a complete submission should add the full $(\delta,p)$ grid for fatigue decisions, alert burden, and benchmark-relative loss estimates.
@@ 6.5 Case Study Visualisations
To demonstrate the practical application and interpretability of our methodology, we apply it to seven synthetic case-study trajectories designed to mimic plausible scenarios of creative performance decay and missingness. The resulting visualisations, shown in Figures 1 through 7, illustrate how the signature-based method identifies change points and classifies performance trends across a variety of fatigue patterns. Each plot displays the daily CTR, the detected change points as vertical dashed lines, and the classified trend for each segment.
@@
Figure 6: Non-Continuous Data (Sharp Drop): Illustrates the detector on a trajectory with significant gaps in the data.
Figure 7: Non-Continuous Data (Gradual Decline): The signature method detects a change point despite irregular sampling.
These visualisations illustrate the algorithm on several fatigue-pattern examples. For a marketing analyst, these plots provide a narrative of a creative’s performance lifecycle, highlighting not just *that* performance has changed, but *when* and in what manner. Figures 6 and 7 provide qualitative examples under missingness; they are illustrative rather than a general robustness guarantee. Window-based baselines can also be applied under missingness, but require an explicit choice of whether windows are defined over calendar days or observed days.
@@ 7 Results
We evaluate detection performance against the operational fatigue onset definition in Section 6.4. The dataset contains 17 creatives over 2025–01–01 to 2025–06–30, of which 6 meet the operational fatigue onset criterion under the main specification ($\delta=0.15$, $p=5$, $w_{s}=7$, $B=21$, $m_{0}=1000$). For each forward-trigger baseline, we define an alert date $\hat{\tau}$ as the first calendar day on which that method’s alert condition is satisfied. If a method requires persistence for three observations, $\hat{\tau}$ is the first day of that qualifying run. For the signature-based approach in this panel experiment, we instead report a retrospective segment-start marker: the start date of the first segment classified as “declining” (with a fallback to the first detected change point). The signature threshold uses the full-series distribution of distances, so this marker can be earlier than the day on which an online detector could have generated an actionable alert. Online deployment requires a past-only calibration rule and should be evaluated separately.
@@
Table 1: Retrospective timing and alert burden under the rule-based onset definition
Method $N_{fat}$ $N_{hit}$ $N_{miss}$ $N_{burden}$ Median signed timing $L$
@@ 7.1 Interpretation: early warning versus alert burden
Table 1 highlights a practical trade-off. The signature-based row marks a declining segment before the rule-based onset for all fatigue creatives, and the moving-average crossover triggers before the rule-based onset for all fatigue creatives under this definition. Both also flag all non-fatigue creatives in this dataset. At this operating point, they therefore provide no creative-level filtering on the panel experiment: every creative is flagged at least once. By contrast, the rolling regression baseline has zero non-fatigue creative-level burden in this table, but misses most fatigue onsets and is, on average, late when it does trigger. Large positive signed timing values should be interpreted cautiously: extremely early markers can reflect a mismatch between a heuristic trigger and the operational onset definition, rather than “better” detection. For the signature-based detector and the moving-average crossover, the current results motivate burden-constrained evaluation and stronger baselines rather than a claim of creative-level discrimination.
7.2 Injected changepoint benchmark
To separate agreement with the operational rule from recovery of known event times, we also run a stochastic benchmark with a deterministic injected changepoint at a known onset time. The benchmark contains 264 simulated creative trajectories of length 180 days: 72 event trajectories (abrupt drop, gradual decline, and volatile decline) and 192 no-event trajectories. The no-event regimes include stable controls, impression-volume shifts, seasonality changes, delivery gaps, targeting-mix shifts, auction-pressure shocks, attribution-reporting volatility, and creative-rotation cycles. Event scenarios have a known onset at day 90. Daily clicks are generated from a Binomial model conditional on impressions, with impression volumes drawn from a noisy weekly seasonal process. CTR is smoothed using the same Beta–Binomial shrinkage logic used in the rule-based panel evaluation. Because the detector is applied to Beta–Binomial smoothed CTR rather than latent CTR, impression-volume shifts can change the observed benchmark series through the shrinkage map even when latent click probability is unchanged. Some no-event regimes deliberately change observed or latent CTR for non-fatigue reasons; these cases test whether a detector over-interprets operational confounding as creative wear-out.
Table 2: Synthetic injected-benchmark diagnostics
Group Series Days Median impr. Impr. IQR CTR p10–p90
Known-onset event 72 180 7358 4102 1.35–2.55%
No-event control 192 180 7195 4862 1.49–2.05%
Table 2 summarises the synthetic benchmark scale. It is included to make the synthetic-only evidence auditable: the simulations cover daily histories with thousands of impressions, low-single-digit observed CTR, and controlled no-event nuisance regimes, including simplified versions of several advertising-specific confounds. They still do not reproduce proprietary platform mechanics, auction systems, or human creative-review decisions. The results below should therefore be interpreted as internal validity and stress-test evidence under stated data-generating processes.
Table 3: Synthetic scenario diagnostics around the injected reference day
Scenario Type Latent CTR pre/post Observed CTR pre/post Impr. pre/post
Stable No event 1.80/1.80% 1.81/1.80% 7040/7076
Impression shift No event 1.80/1.80% 1.80/1.80% 8466/3842
Seasonality shift No event 1.80/1.80% 1.80/1.80% 7704/7704
Delivery gaps No event 1.80/1.80% 1.81/1.79% 7349/7418
Targeting mix No event 1.80/1.45% 1.80/1.45% 8411/8550
Auction pressure No event 1.80/1.80% 1.80/1.81% 7591/5460
Attribution volatility No event 1.80/1.80% 1.76/1.84% 8355/8448
Creative rotation No event 1.81/1.79% 1.81/1.80% 7956/8383
Abrupt drop Event 2.40/1.50% 2.39/1.50% 7390/7528
Gradual decline Event 2.40/1.66% 2.39/1.65% 8201/8328
Volatile decline Event 2.40/1.50% 2.40/1.50% 7561/7611
Table 3 gives a compact audit of the scenario mechanisms by comparing median pre/post quantities around the injected reference day. The event rows show deliberate post-onset degradation in latent CTR. The no-event nuisance rows are not all stationary: targeting mix and attribution volatility can move observed or latent CTR for reasons other than creative fatigue, while auction pressure, impression shifts, and delivery gaps primarily affect exposure. These rows are intentionally difficult controls; a useful screen should surface their risk separately rather than treating every post-reference deterioration as wear-out.
For this benchmark, an event hit is defined as the first actionable alert falling in $\tau,\tau+45]$, where $\tau$ is the injected onset. For adjacent-window online methods, the scored boundary is between the left and right windows, but the actionable alert date is the final day of the right window because those observations are required to compute the statistic. Alerts before $\tau$ or after $\tau+45$ are counted as invalid event alerts, and any alert on no-event controls is counted as a no-event false alert. We evaluate the original shape-only signature detector separately from level-aware variants. The shape-only variant uses independent per-window min–max scaling and no directional mean-drop gate. The level-aware variants use calibration-standardised CTR so that absolute deterioration is retained, and require the post-change window mean to be at least two calibration standard deviations below the pre-change window mean. This separation makes clear which evidence supports the original geometric representation and which evidence supports the fatigue-specific level-aware extension.
Table 4: Injected changepoint benchmark with nuisance controls
Method Event hit (72) Median delay Invalid alert (72) No-event false (192) Series
Signature shape-only 26% 15 72% 99% 72/192
Level-aware sig. retros. 29% 6 71% 25% 72/192
Level-aware signature 94% 6 6% 31% 72/192
Level+exposure signature 96% 6 4% 32% 72/192
Window features 94% 6 4% 33% 72/192
Kernel window 93% 11 7% 47% 72/192
GLR mean-drop 90% 6 10% 49% 72/192
MOSUM mean-drop 50% 6 50% 93% 72/192
CUSUM drop 78% 5 22% 57% 72/192
Offline full-series ref. 100% 4 0% 12% 72/192
Table 4 is deliberately conservative: pre-onset alerts are not credited as hits, and adjacent-window online methods are scored at the day when the right window has become observable. Under this scoring rule, the shape-only signature detector has limited value for a fatigue DGP dominated by level deterioration, while the primary level-aware online signature variant recovers most known events with lower no-event burden than several simpler online comparators. The hand-crafted window-feature baseline remains a close competitor, and the Gaussian-kernel and GLR-style baselines also recover many events but with higher nuisance burden. This is important: the benchmark supports the level-aware adjacent-window workflow, but it does not show that signatures dominate all non-signature window summaries. The MOSUM and GLR rows are intentionally strong online level-change comparators; their results help distinguish the value of the signature representation from the value of simply detecting adjacent-window mean deterioration. The offline full-series reference uses future observations and is not an operational online competitor; its row should be read as an oracle-style upper reference for this simplified one-change design. It performs best on this benchmark, which is expected because the data-generating process matches its assumptions more closely than it matches the broader dynamic-pattern objective of path signatures. This result narrows the claim: signatures are a useful screening statistic for some heterogeneous trajectories, but the present benchmark does not establish superiority over all strong feature-based changepoint screens.
Table 5: Operating points for online detectors under matched scoring
Method / setting Event hit Median delay Invalid event alert No-event false alert
Signature shape-only ($k=1.0$) 8% 4 92% 100%
Signature shape-only ($k=1.5$) 26% 15 72% 99%
Signature shape-only ($k=2.5$) 18% 22 28% 56%
Level-aware signature ($k=1.0$) 94% 6 6% 35%
Level-aware signature ($k=1.5$) 94% 6 6% 31%
Level-aware signature ($k=2.5$) 94% 6 4% 26%
Level+exposure signature ($k=1.0$) 96% 6 4% 36%
Level+exposure signature ($k=1.5$) 96% 6 4% 32%
Level+exposure signature ($k=2.5$) 99% 6 1% 24%
Window features ($k=1.0$) 92% 6 7% 39%
Window features ($k=1.5$) 94% 6 4% 33%
Window features ($k=2.5$) 83% 5 8% 24%
Kernel window ($k=1.0$) 90% 10 10% 49%
Kernel window ($k=1.5$) 93% 11 7% 47%
Kernel window ($k=2.5$) 53% 13 3% 32%
GLR mean-drop ($k=1.0$) 90% 6 10% 49%
GLR mean-drop ($k=1.5$) 90% 6 10% 49%
GLR mean-drop ($k=2.5$) 90% 6 10% 49%
MOSUM mean-drop ($k=1.0$) 29% 5 71% 100%
MOSUM mean-drop ($k=1.5$) 50% 6 50% 93%
MOSUM mean-drop ($k=2.5$) 90% 6 10% 48%
[…]
Table 5 reports selected operating points under the same post-onset hit definition. The table makes the alert-budget trade-off explicit. For the level-aware variants, increasing the signature threshold lowers no-event false alerts but does not materially improve event recovery in this simulation. The shape-only rows are not monotone in event-hit rate because changing the threshold can move alerts from invalid pre-onset times into the post-onset scoring window; this reinforces that shape-only signatures are not the recommended fatigue screen. The kernel-window rows show that a generic distribution-distance comparator can recover many events, but its operating points are either slower and higher burden or substantially lower recall. The GLR-style mean-drop rows behave like a strong level-change screen with high recovery but higher nuisance burden than the level-aware signature rows. These operating curves are more informative than a single headline setting and support treating the method as a candidate screen whose threshold must be tuned to an explicit alert budget.
Table 6: False-alert rates by no-event nuisance scenario for selected online methods
Nuisance regime Shape Level Exp. Feat. Kernel GLR CUSUM
Attribution volatility 100% 88% 75% 83% 92% 92% 100%
Auction pressure 100% 4% 8% 4% 38% 38% 42%
Creative rotation 100% 0% 0% 0% 0% 0% 4%
Delivery gaps 96% 21% 25% 29% 50% 62% 67%
Impression shift 100% 29% 25% 29% 42% 50% 50%
Seasonality shift 100% 8% 17% 12% 25% 25% 46%
Stable 100% 4% 12% 12% 29% 29% 50%
Targeting mix 100% 96% 96% 92% 100% 100% 100%
Table 6 reports false-alert rates separately by no-event nuisance scenario. The most difficult controls differ by method, but the broad pattern is that shape-only signatures are highly sensitive to non-fatigue trajectory changes, while level-aware variants reduce but do not eliminate nuisance alerts. This supports the ReviewBot concern that signatures are not, by themselves, a causal fatigue model. The kernel-window and GLR comparators are also sensitive to several nuisance regimes, especially targeting mix and attribution volatility, which reinforces that generic window dissimilarity or level-drop evidence is not enough without domain constraints. The exposure-aware signature variant adds useful information in some volume-driven regimes, but it remains sensitive to targeting, auction, attribution, and rotation mechanisms that are deliberately outside the fatigue event definition.
Table 7: Signature-depth ablation for the past-only detector
Depth Event hit Median delay No-event false alert
2 94% 6 33%
3 94% 6 31%
4 96% 6 32%
Table 7 shows that increasing depth from 2 to 4 does not materially improve this benchmark. This is useful negative evidence: for the simulated fatigue mechanisms considered here, most of the recoverable signal is already available at low signature depth. More complex creative trajectories may require higher-depth or log-signature features, but the current evidence does not depend on depth 4.
Table 8: Component ablation for the past-only level-aware detector
Variant Event hit Median delay Invalid alert No-event false alert
Shape-only signature 26% 15 72% 99%
Directional gate only 90% 6 10% 49%
Level-aware sig., no gate 33% 2 67% 96%
Level-aware sig. + gate 94% 6 6% 31%
Level+exposure sig. + gate 96% 6 4% 32%
Table 8 separates the main ingredients of the level-aware detector. The directional gate alone is a simple mean-drop screen; the level-aware signature without the gate measures unusual adjacent-window path changes but does not enforce fatigue direction. The combined level-aware signature plus gate is the primary candidate rule, while the exposure-aware row adds standardised log-impressions as an additional path coordinate. In this benchmark, the gate supplies most of the event recovery, while adding the signature threshold reduces no-event burden relative to the gate-only rule. Exposure augmentation does not eliminate nuisance false alerts. The ablation therefore supports a restrained interpretation: signatures contribute a structured trajectory-change score, but the domain-specific deterioration gate carries important practical content.
Table 9: Benchmark-relative impact proxy after detector alert
Method Median missed clicks Post-onset loss flagged
Signature shape-only 0 100%
Level-aware sig. retros. 0 100%
Level-aware signature 227 95%
Level+exposure signature 231 95%
Window features 231 95%
Kernel window 334 92%
GLR mean-drop 214 96%
MOSUM mean-drop 0 100%
CUSUM drop 136 97%
Offline full-series ref. 195 96%
Table 9 connects alert timing to the benchmark-relative lost-clicks proxy from Section 5. The table reports the median lost clicks that occur after the injected onset but before the first detector alert, along with the median share of post-onset loss covered by the alert timing. The summary includes all event trajectories: pre-onset alerts have zero post-onset missed loss, while no-alert trajectories have all post-onset loss counted as missed. A high captured-loss value can therefore be produced by over-early or over-frequent alerts, so the invalid-alert and no-event burden columns in Table 4 must be read alongside this proxy table. These are not causal financial estimates; they are a transparent way to compare prioritisation consequences under different alert rules.
Table 10: Cost-sensitive operating-point selection on injected benchmark
Method / selected setting Event hit No-event false alert Alert rate Captured clicks Net clicks
Level-aware signature ($k=2.0$) 96% 28% 48% 1389 1318
Level+exposure signature ($k=2.5$) 99% 24% 45% 1384 1317
[…]
Window features ($k=1.5$) 94% 33% 51% 1382 1306
GLR mean-drop ($k=3.0$) 90% 48% 62% 1394 1301
MOSUM mean-drop ($k=2.5$) 90% 48% 62% 1391 1297
Signature shape-only ($k=1.0$) 8% 100% 100% 1431 1281
Kernel window ($k=1.0$) 90% 49% 63% 1365 1271
Table 10 adds a stylised operating-point exercise. For each online method family, it selects the threshold with the highest mean net proxy clicks per simulated series after charging a review cost of 150 clicks for every alert. The objective is
[…]
where alert rate is the fraction of simulated series producing at least one review. Net clicks are computed from unrounded alert rates and captured-click values; percentages in the table are rounded for display. The units are click-equivalent proxies rather than financial estimates. This table gives a concrete way to choose a threshold when analyst attention is costly: a method with high event recovery can still be unattractive if it creates many low-value reviews.
Table 11: Train-selected thresholds evaluated on held-out synthetic replicates
Method / frozen setting Train net Holdout hit Holdout false alert Alert rate Holdout net
[…]
GLR mean-drop ($k=3.0$) 1353 86% 49% 63% 1248
Level+exposure signature ($k=3.0$) 1379 94% 25% 45% 1245
MOSUM mean-drop ($k=2.5$) 1350 89% 52% 65% 1244
Window features ($k=1.5$) 1369 89% 40% 55% 1242
Level-aware signature ($k=3.0$) 1378 89% 26% 45% 1233
Kernel window ($k=1.0$) 1330 86% 51% 64% 1212
Table 11 turns the same idea into a train/held-out protocol. Thresholds are selected on the first 12 replicates of each synthetic scenario using the net-click objective above, then frozen and evaluated on the remaining 12 replicates. The level-aware signature setting remains competitive with the feature, kernel, and GLR baselines under this synthetic split and produces a lower no-event false-alert rate than the CUSUM, GLR, and kernel settings selected by the same objective. Because the split is over replicates from the same scenario families, this table should be read as a leakage-control check, not as evidence that the synthetic nuisance library is complete.
Table 12: Fixed-capacity alert triage by alert-time opportunity proxy
Method Capacity Reviewed Precision Post-onset loss captured
Level-aware signature 10 10 100% 18%
Level-aware signature 20 20 85% 31%
Level+exposure signature 10 10 100% 18%
Level+exposure signature 20 20 90% 33%
Window features 10 10 100% 18%
Window features 20 20 85% 31%
Kernel window 10 10 100% 19%
Kernel window 20 20 80% 27%
GLR mean-drop 10 10 100% 18%
GLR mean-drop 20 20 80% 29%
CUSUM drop 10 10 70% 13%
CUSUM drop 20 20 65% 23%
Table 12 reframes the same alerts as a finite analyst-capacity problem. Within each detector’s alerts, candidate reviews are ranked by the alert-time lost-click proxy computed from the calibration benchmark and the current posterior CTR. The table reports the top 10 and top 20 reviews for selected online methods. Precision is the share of reviewed alerts that fall inside the post-onset hit window, while loss captured is the share of total injected post-onset loss associated with those reviewed alerts. This is a triage diagnostic rather than a causal value estimate, but it makes the practical capacity trade-off more explicit than detector hit rates alone.
Table 13: Sensitivity of opportunity ranking to benchmark definition
Benchmark rule Median lost clicks Rank corr. Top-10 overlap
Calibration mean 5097 1.00 100%
Calibration 75th pct. 5687 1.00 90%
Pre-onset recent best 6306 0.98 90%
Table 13 checks whether the opportunity ranking is stable to reasonable benchmark definitions. In this injected benchmark, the rank correlation with the calibration-mean benchmark remains high and the top-10 overlap is at least 90%, although the median lost-click magnitude rises when the benchmark uses a higher pre-onset reference. This supports the proxy as a ranking aid in the synthetic setting, while showing that absolute loss magnitudes remain benchmark-dependent.
Table 14: Worked toy signature distances for simple CTR paths
Contrast Same endpoints Depth 1 Depth 2 Depth 3
monotone vs zig-zag Yes 0.00000 0.00000 0.00055
monotone vs recover/fall Yes 0.00000 0.00707 0.00791
monotone vs flat No 0.01000 0.01225 0.01258
Table 14 gives a small representation sanity check. The monotone and zig-zag paths have the same start and end CTR, so their depth-1 displacement distance is zero. Higher-depth terms can distinguish some equal-endpoint path shapes, as seen in the depth-3 distance for the zig-zag contrast and the depth-2/3 distances for the recover-then-fall contrast. This toy example illustrates the geometric contribution of signatures, but it is not by itself evidence of superior detector performance under noisy advertising data.
Table 15: Worked online level-aware detector trace
Case Boundary Available Raw CTR L/R $\widetilde{CTR}$ L/R $D$ $\theta$ $G$ / decision
Abrupt drop 84 97 2.31/1.88% 2.30/2.14% 161.98 34.77 2.02 / alert
Targeting mix 85 98 1.84/1.59% 1.81/1.69% 82.04 51.36 2.16 / false alert
Stable 100 1.79/1.78% 1.80/1.76% 7.71 7.36 0.88 / no alert
Table 15 walks through the primary online detector on three benchmark cases. The boundary column is the adjacent-window split being scored; the available column is the first day on which the alert can be issued after observing the right window. The abrupt-drop row shows the intended mechanism: raw and smoothed CTR deteriorate, the signature distance exceeds its past-only threshold, and the directional gate is satisfied. The targeting-mix row is deliberately included as a failure mode: the same conditions can be met by a non-fatigue regime shift, so the method should be treated as a screen for analyst review rather than a causal fatigue classifier. The stable row shows a no-alert case where the displayed boundary is the closest eligible score to a joint threshold crossing.
Table 16: Runtime smoke benchmark for past-only signature detection
Series Depth Seconds Projected sec. / 1000 series
25 3 0.54 21.5
25 4 0.64 25.5
100 3 2.14 21.4
100 4 2.51 25.1
Finally, Table 16 provides a small runtime smoke benchmark for past-only signature detection on replicated synthetic series using full-path roughpy signatures. The results are intended to validate order-of-growth plausibility, not to replace an end-to-end system benchmark. They suggest that fixed-depth portfolio screening is computationally feasible at the scale of thousands of daily creative trajectories on the test machine, while a deployed system would still need timing that includes data loading, report generation, storage, and alert delivery.
7.3 Synthetic-only evidence boundary
No client campaign histories are used in this paper. This avoids confidentiality risk and keeps the results fully reproducible, but it also limits the claims. The experiments can show whether the proposed detectors behave sensibly under transparent synthetic data-generating processes with known onsets and nuisance controls. They cannot establish that the detector will identify fatigue in a particular production portfolio, that the synthetic nuisance set spans all operational confounders, or that the benchmark-relative opportunity proxy will match analyst-reviewed business value. For this reason, the contribution is framed as a synthetic evaluation of a screening framework rather than as a production-calibrated fatigue detector.
Under a synthetic-only evidence standard, the main quality controls are specification, coverage, and reproducibility. The data-generating mechanisms, onset scoring rule, no-event scenarios, detector variants, and generated tables are all included in the repository under presentations/paper/experiments. Future synthetic extensions should deepen the nuisance library further to include platform learning phases, intervention feedback loops, multi-stage creative deterioration, and cross-campaign dependence while preserving the same aggregate reporting discipline.